In any triangle, the sum of the measures of the angles is 180Degrees. In triangle ABC, Angle A is 4 times as large as Angle B. Angle C measures 20degrees less than Angle B. Find the measure of each angle.
I know he answer are the following but I need help figuring out how to write the equation and then help to solve A=133.32 B=33.33 C=13.33
Please help!

Answer by scott8148(6628)

"Angle A is 4 times as large as Angle B" ___ A = 4B

"Angle C measures 20degrees less than Angle B" ___ C = B - 20

A + B + C = 180

4B + B + B - 20 = 180 ___ 6B - 20 = 180 ___ 6B = 200 ___ B = 200/6 = 100/3

substituting ___ A = 4 (200/6) = 800/6 = 400/3

substituting ___ C = 200/6 - 120/6 = 80/6 = 40/3